# FinLoRA: Finetuning Quantized Financial Large Language Models Using Low-Rank Adaptation

### v1

@@ Abstract
Finetuned large language models (LLMs) have shown remarkable performance in financial tasks, such as sentiment analysis and information retrieval. Due to privacy concerns, finetuning and deploying Financial LLMs (FinLLMs) locally are crucial for institutions. However, finetuning FinLLMs poses challenges including GPU memory constraints and long input sequences. In this paper, we employ quantized low-rank adaptation (QLoRA) to finetune FinLLMs, which leverage low-rank matrix decomposition and quantization techniques to significantly reduce computational requirements while maintaining high model performance. We also employ data and pipeline parallelism to enable local finetuning using cost-effective, widely accessible GPUs. Experiments on financial datasets demonstrate that our method achieves substantial improvements in accuracy, GPU memory usage, and time efficiency, underscoring the potential of low-rank methods for scalable and resource-efficient LLM finetuning.
@@ Introduction
Large language models (LLMs) have demonstrated exceptional capabilities in various applications, such as finance (Liu et al. 2023a, 2024b), healthcare (Wang et al. 2024; Chen et al. 2024b), scientific discovery (Lu et al. 2022b; Chen et al. 2021, 2024a; Fu et al. 2024), etc. Finetuning using low-rank structures of these models to domain-specific datasets further enhances their performance and improves their applicability to specialized tasks. In the financial domain, finetuned LLMs demonstrate substantial potential for tasks such as sentiment analysis, named-entity recognition, and knowledge extraction from financial documents.
FinGPT (Liu et al. 2023a, 2024b, 2024a) applied low-rank adaptation techniques for finetuning quantized LLMs in financial contexts, which displayed noticeable improvement over the base model, while having substantial memory reduction and training speedup. XBRL agent (Han et al. 2024) evaluated the potential of LLMs’ capabilities on analyzing XBRL reports. The use of Retrieval-Augmented Generation (RAG) and tools-calling techniques on XBRL-related tasks and demonstrated significant improvement in task accuracy.
Due to sensitive data and regulatory constraints, finetuning and inference of LLMs within local environments remain critical requirements for financial institutions. Furthermore, the ability to create personalized and customized LLMs, finetuned for specific tasks, is essential for maximizing the utility of these models in financial applications.
Challenges Existing FinLLMs face the following challenges:
  • Inefficient finetuning: Financial tasks are often complex and require precise adaptation to large domain-specific data, which can involve extensive parameter updates and prolonged training times.
  • Resource-constrained local devices: Finetuning and deploying large language models on resource-constrained local devices, such as consumer GPUs, can be challenging due to their memory and computational limitations.
Building upon prior research, we demonstrate that state-of-the-art LLMs can be finetuned for diverse financial tasks locally and cost-effectively using widely accessible GPUs, achieving notable improvements over baseline models. Our main contributions can be summarized as follows:
  1. 1.
    We employ Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al. 2023) to alleviate memory requirements and allow more efficient finetuning. Using the low-rank structure reduces the number of trainable parameters required for finetuning, and quantization compresses the model size, further limiting GPU memory consumption.
  2. 2.
@@
Low-rank adaptation (LoRA) (Hu et al., 2021) is a parameter-efficient finetuning method that preserves the pretrained transformer model weights and introduces a smaller set of trainable weights, which are expressed using low-rank decomposition.
In LoRA, the update weights are assumed to follow the low-rank decompositions $\Delta\bm{W}=\bm{B}\bm{A}$, where $\bm{A}\in\mathbb{R}^{r\times n}$ and $\bm{B}\in\mathbb{R}^{n\times r}$ are trainable parameters, $\bm{W}_{0}\in\mathbb{R}^{n\times n}$ is the pretrained weights. Note that $n$ can be large, e.g., $4096$ and the rank $r\ll n$, say $r=4,8$, or $16$. As an example, setting $n=4096$ and $r=8$, $\bm{W}_{0}$ has about $16$ million parameters, while $\bm{A}$ and $\bm{B}$ together have $65536$ parameters, which is about 0.039% the size of $\bm{W}_{0}$.
During the finetuning stage, the forward pass is:
@@
where $\bm{W}_{0}$ denotes the pretrained weights.
During the inference stage, $\bm{A}$ and $\bm{B}$ will be added back to $\bm{W}_{0}$, resulting in matrix $\bm{W}$ with the size of $\bm{W}_{0}$, such that
$\bm{W}=\bm{W}_{0}+\bm{B}\bm{A},$
$\bm{y}=\bm{W}\bm{x}.$
Therefore, it does not introduce additional costs to the inference process.
@@
Table 1: The maximum GPU memory used during inference.
@@
To accelerate the finetuning process and leverage the computational power of multiple GPUs, we employed Distributed Data Parallel (DDP), which distributes the training data across multiple GPUs. DDP launches one process per GPU, where each process gets its own copy of the model and optimizer. Each process then receives different inputs, and the gradients are computed and synchronized across all GPUs to accelerate training. DDP offers significant speedup when multiple GPUs are available (Li et al. 2020).
We also opted to use Brain Floating Point (BF16) during finetuning, BF16 offers range of values the same as FP32 and easy conversion to/from FP32. Studies showed that BF16 can achieve similar results as FP32 while having significant speedup and memory savings (Kalamkar et al. 2019).
We used 0/1 Adam optimizer (Lu et al. 2022a), a modified version of the Adam optimizer that linearizes each Adam step and allows utilizing 1-bit compression for faster convergence speed, while offering reduced data volume and higher training throughput.
@@
Inference on larger models like Llama 3.1 70B requires substantial GPU memory usage, particularly when using higher precision like 8-bit or 16-bit. We employ pipeline parallelism, where the model is partitioned at the layer level and distributed across multiple GPUs; each GPU process computes different micro-batches with different parts of the model concurrently (Liu et al. 2024a).
Table 1 illustrates GPU memory usage achieved through quantization and pipeline parallelism during inference with a batch size of 1. For Llama 3.1-8B, memory usage decreases by 43% with 8-bit quantization and 63% with 4-bit quantization compared to the original 16-bit. Similarly, for Llama 3.1-70B, the memory requirement reduces from 131.5 GB (16-bit, requiring 3 GPUs) to 68.5 GB (8-bit, 2 GPUs) and further to 37.8 GB (4-bit, 1 GPU). These reductions demonstrate the practical benefits of quantization in enabling resource-efficient inference for large-scale models.
@@
Hardware Configurations. The experiments were conducted on a server equipped with four 16-core AMD EPYC 7313 CPUs, 1 TB of RAM, and four NVIDIA RTX A6000 GPUs, each featuring 48 GB of dedicated GPU memory.
@@
For classification tasks, our study focuses on three financial language processing tasks: Sentiment Analysis (SA), Named Entity Recognition (NER), and news headline classification.
@@
For Question-Answering (QA) tasks, we focus on eXtensible Business Reporting Language (XBRL) (Saeedi, Richards, and Smith 2007) data extraction. XBRL is a standardized format designed for the exchange of financial information. Although XBRL documents are based on structured XML (eXtensible Markup Language), their inherent complexity presents challenges that can be addressed using the capabilities of LLMs to extract key information, thereby facilitating financial reporting and analysis (Han et al. 2024). In this study, we aim to finetune LLMs to accurately extract both numerical and textual information from XBRL files.
Base Models
@@
Dataset Name Type Train/Test Examples
FPB Classification 1,200 / 3,600
FiQA SA Classification 961 / 150
TFNS Classification 9,540 / 2,390
NWGI Classification 16,200 / 4,050
Headline Classification 82,200 / 20,500
NER Classification 13,500 / 3,500
XBRL Tags QA 375 / 164
XBRL Values QA 846 / 154
Table 3: Accuracy on classification and XBRL extraction tasks.
Model Classification Datasets XBRL Extraction
FPB FIQA TFNS NWGI NER Headline Tags Values
Llama-3.1-8B-Instruct (base) 0.6873 0.4655 0.6997 0.4658 0.4889 0.4534 0.7937 0.5526
Llama-3.1-70B-Instruct (base) 0.7450 0.4727 0.6842 0.7993 0.4628 0.7168 0.8902 0.8766
Llama-3.1-8B-Instruct-4bits-r4 0.8630 0.7309 0.8827 0.8095 0.9663 0.8803 0.9500 0.9605
Llama-3.1-8B-Instruct-8bits-r8 0.8284 0.8036 0.8405 0.8396 0.9805 0.8466 0.9437 0.9736
Llama-3.1-70B-Instruct-4bits-r4 - - - - 0.9888 - - -
Table 4: Weighted F1 score on all classification tasks.
Model Classification Datasets
FPB FIQA TFNS NWGI NER Headline
Llama-3.1-8B-Instruct (Base) 0.6768 0.5571 0.6834 0.4117 0.5686 0.5576
Llama-3.1-70B-Instruct (Base) 0.7363 0.5645 0.6864 0.7993 0.4539 0.7294
Llama-3.1-8B-Instruct-4bits-r4 0.8600 0.7811 0.8824 0.8029 0.9664 0.8864
Llama-3.1-8B-Instruct-8bits-r8 0.8302 0.8177 0.8436 0.8492 0.9806 0.8520
Llama-3.1-70B-Instruct-4bits-r4 - - - - 0.9887 -
@@
Llama-3.1-8B-Instruct-4bits-r4 16 $\times$ 4 83.6 0.77 $\times$ 4 4.5 0.1
Llama-3.1-8B-Instruct-8bits-r8 16 $\times$ 4 96.7 0.90 $\times$ 4 9.0 0.1
Llama-3.1-70B-Instruct-4bits-r4 4 $\times$ 4 184.3 3.50 $\times$ 4 21.3 0.9
Table 6: Finetuning and inference performance on XBRL.
@@
Batch size GPU memory (GB) GPU hours Adapter size (MB) Time per Example (s)
Llama-3.1-8B-Instruct-4bits-r4 2 $\times$ 4 139.2 0.44 $\times$ 4 4.5 1.9
Llama-3.1-8B-Instruct-8bits-r8 2 $\times$ 4 152.2 0.48 $\times$ 4 9.0 1.9
@@
The following datasets all consist of input texts and sentiment labels such as "neutral", "positive", or "negative".
@@
  • News with GPT instruction (NWGI) (Liu et al. 2023b) comprises samples with seven labels ranging from “strong negative” to “strong positive”. We map the seven labels back to three for simplicity and consistancy with other SA dataset.
@@
The Headline dataset (Sinha and Khandait 2020) classifies headlines based on various questions into two classes: "yes" and "no".
@@
The NER dataset (Salinas Alvarado, Verspoor, and Baldwin 2015) annotates one entity per sentence, categorized into one of three classes: "location", "person", and "organization"
XBRL
The XBRL dataset (Han et al. 2024) comprises questions and answers derived from XBRL filings from 2019 to 2023 for Dow Jones 30 companies. Each example includes a question, a text segment from an XBRL file containing the answer, and the ground truth generated using an XBRL file extraction library. From this dataset, we selected the following two tasks:
  • XBRL tag extraction: This task involves extracting a specific XBRL tag from a large XBRL raw text segment given a natural language description of the tag.
  • XBRL value extraction: This task focuses on extracting a numeric value from the raw XBRL text segment given a natural language description of the value.
@@
  • Classification Tasks: For classification tasks, including sentiment analysis, headline classification, and named entity recognition, single-task fine-tuning was employed.
  • XBRL Question Answering: For XBRL question-answering tasks, including tag extraction and value extraction, multi-task fine-tuning was adopted.
@@
  • Classification Tasks:
@@
Tables 3 and 4 summarize the accuracy and weighted F1 scores under different finetuning configurations. Table 5 and 6 displays resource usage and inference performance for NER and XBRL QA. The finetuned Llama 3.1 8B demonstrates noticeable improvements in accuracy compared to its base model and even surpasses the results of the Llama 3.1 70B base model.
@@
This study demonstrates the effectiveness of Quantized LoRA (QLoRA) for finetuning large language models (LLMs) on a range of financial tasks, including sentiment analysis, named entity recognition, news headline analysis, and XBRL data extraction. We finetuned both Llama 3.1 8B and 70B models, achieving up to 48% improvements in accuracy compared to the base models on average across all tasks. Notably, these performance gains can be achieved with only four GPUs and less than 20 hours of training times per task, making local finetuning and deployment of customized models a feasible option for financial institutions.
In future work, we plan to explore multi-task finetuning in classification tasks and expand our investigation of XBRL-related tasks, including formula calculations. This will enable FinLLMs to perform more complex analysis and reasoning tasks, further increasing their utility in the financial domain.

### v2

@@ Abstract
Finetuned large language models (LLMs) have shown remarkable performance in financial tasks, such as sentiment analysis and information retrieval. Due to privacy concerns, finetuning and deploying financial LLMs (FinLLMs) locally are crucial for institutions and individuals. In this paper, we employ quantized low-rank adaptation (QLoRA) to finetune FinLLMs, which leverage low-rank structure and quantization technique to significantly reduce computational requirements while maintaining model performance. We also employ data and pipeline parallelism to enable local finetuning on commodity GPUs. Experiments on financial datasets validate the efficacy of our approach in yielding notable improvements over the base models.
@@ Introduction
Large language models (LLMs) have demonstrated exceptional capabilities in various applications, such as finance (Liu et al. 2023, 2024b), healthcare (Wang et al. 2024; Chen et al. 2024c, a), scientific discovery (Lu et al. 2022b; Chen et al. 2021, 2024b; Fu et al. 2024), etc. Finetuning using low-rank structures of these models to domain-specific datasets further enhances their performance and improves their applicability to specialized tasks. In the financial domain, finetuned LLMs demonstrate substantial potential for tasks such as sentiment analysis, named entity recognition (NER), and knowledge extraction from financial documents.
FinGPT (Liu et al. 2023, 2024b, 2024a; Tian et al. 2024) applied low-rank adaptation techniques for finetuning quantized LLMs in financial contexts, which displayed noticeable improvement over the base model, while having substantial memory reduction and training speedup. XBRL agent (Han et al. 2024) evaluated the potential of LLM’s capabilities in analyzing XBRL reports. The use of Retrieval-Augmented Generation (RAG) and tools-calling techniques on XBRL-related tasks and demonstrated significant improvement in task accuracy.
Due to sensitive data and regulatory constraints, finetuning and inference of LLMs within local environments remain critical requirements for financial institutions. Furthermore, the ability to create personalized and customized LLMs, finetuned for specific tasks, is essential for maximizing the utility of these models in financial applications, such as the FinGPT search agents (Tian et al. 2024).
Building upon prior research (Liu et al. 2024a), we demonstrate that state-of-the-art LLMs can be finetuned for diverse financial tasks locally and cost-effectively using widely accessible GPUs, achieving notable improvements over baseline models. Our main contributions can be summarized as follows:
  • We employ Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al. 2023) to alleviate GPU memory requirements and achieve more efficient finetuning. Using the low-rank structure reduces the number of trainable parameters required for finetuning, and quantization compresses the model size, further reducing GPU memory consumption.
@@
Our codes are available at Github [^1] FinLoRA: https://github.com/YangletLiu/FinLoRA
@@
Low-rank adaptation (LoRA) (Hu et al. 2021) is a parameter-efficient finetuning method that incorporates a smaller set of trainable weights, such that $\bm{W}=\bm{W}_{0}+\Delta\bm{W}$. Let $\bm{W}_{0}\in\mathbb{R}^{n\times n}$ denote the pretrained weights and $\Delta\bm{W}=\bm{B}\bm{A}$ denote the update weights, where $\bm{A}\in\mathbb{R}^{r\times n}$ and $\bm{B}\in\mathbb{R}^{n\times r}$ are trainable parameters. Note that $n$ can be large, e.g., $4,096$ and the rank $r\ll n$, say $r=4,8$, or $16$. As an example, setting $n=4,096$, and $r=8$, then $\bm{W}_{0}$ has approximately $16$ million parameters, while $\bm{A}$ and $\bm{B}$ together have $65,536$ parameters, which is approximately only 0.039% the size of $\bm{W}_{0}$.
During the fine-tuning stage, the forward pass can be expressed as:
@@
where $\bm{W}_{0}$ denotes the pre-trained weights.
During the inference stage, we do not add $\bm{A}$ and $\bm{B}$ back to $\bm{W}_{0}$, and we perform
$\bm{y}=\bm{W}_{0}\bm{x}+\bm{B}\bm{A}\bm{x}.$
This is different from (Hu et al. 2021), because we will explore the Mixture of Experts approach that trains multiple LoRA adapters. Therefore, it introduces a small amount of additional costs to the inference process.
@@
Table 1 illustrates GPU memory usage with QLoRA during finetuning with batch size and rank of 8 and inference with a batch size of 1. The reductions in GPU memory with quantization displayed practical benefits of resource-efficient finetuning and inference for large-scale models.
Table 1: GPU memory usage during finetuning (blue, batch size and rank of 8) and inference (black, batch size of 1).
@@
To accelerate the finetuning process and leverage the computational power of multiple GPUs, we employed Distributed Data Parallel (DDP), which distributes the training data across GPUs. DDP launches one process per GPU, where each process gets its own copy of the model and optimizer. Each process then receives different inputs, and the gradients are computed and synchronized across all GPUs to accelerate training. DDP provides a substantial speedup when multiple GPUs are available (Li et al. 2020).
We also opted to use Brain Floating Point (BF16) during finetuning. BF16 offers the same range of values as FP32 an‘d easy conversion to/from FP32. Studies showed that BF16 can achieve similar results as FP32 while having significant speedup and memory savings (Kalamkar et al. 2019).
We used 0/1 Adam optimizer (Lu et al. 2022a), a modified version of the Adam optimizer that linearize each Adam step and allows utilizing 1-bit compression for faster convergence speed, while offering reduced data volume and higher training throughput.
@@
Inference on large-scale models such as Llama 3.1 70B demands considerable GPU memory resources, particularly when using higher precision like 8-bit or 16-bit. We employ pipeline parallelism, where the model is partitioned at the layer level and distributed across multiple GPUs; each GPU process computes different micro-batches with different parts of the model concurrently (Liu et al. 2024a).
@@
The experiments were conducted on a server equipped with four 16-core AMD EPYC 7313 CPUs, 1 TB of RAM, and four NVIDIA RTX A6000 GPUs, each featuring 48 GB of dedicated GPU memory.
@@
For general financial tasks, our study focuses on three financial language processing tasks: Sentiment Analysis (SA), Named Entity Recognition (NER), and news headline classification.
@@
For eXtensible Business Reporting Language (XBRL) (Saeedi, Richards, and Smith 2007) tasks we focus on tagging and extraction. XBRL is a standardized format designed for the exchange of financial information. Although XBRL documents are based on structured XML (eXtensible Markup Language), their inherent complexity presents challenges that can be addressed using the capabilities of LLMs, thereby facilitating financial reporting and analysis (Han et al. 2024).
@@
Dataset Name Type Train/Test Samples
FPB Sentiment Analysis 1.2K / 3.6K
FiQA SA Sentiment Analysis 961 / 150
TFNS Sentiment Analysis 9.5K / 2.4K
NWGI Sentiment Analysis 16.2K / 4.1K
Headline Headline Analysis 82.2K / 20.5K
NER NER 13.5K / 3.5K
FiNER XBRL Tagging 900K / 100K
Tags XBRL Extraction 300 / 150
Values XBRL Extraction 1K / 150
Table 3: Performance on Classification and XBRL Extraction Tasks: Accuracy (blue) and F1 Score (black).
Model Classification Datasets XBRL Extraction XBRL Tagging
FPB FIQA TFNS NWGI NER Headline Tags Values FiNER
Llama-3.1-8B (base) 68.73% 46.55% 69.97% 46.58% 48.89% 45.34% 79.37% 55.26% 2.85%
0.6768 0.5571 0.6834 0.4117 0.5686 0.5576 - - -
Llama-3.1-70B (base) 74.50% 47.27% 68.42% 79.93% 46.28% 71.68% 89.02% 87.66% 9.41%
0.7363 0.5645 0.6864 0.7993 0.4539 0.7294 - - -
Llama-3.1-8B-4bits-r4 86.30% 73.09% 88.27% 80.95% 96.63% 88.03% 95.00% 96.05% 70.45%
0.8600 0.7811 0.8824 0.8029 0.9664 0.8864 - - -
Llama-3.1-8B-8bits-r8 82.84% 80.36% 84.05% 83.96% 98.05% 84.66% 94.37% 97.36% 75.21%
0.8302 0.8177 0.8436 0.8492 0.9806 0.8520 - - -
Llama-3.1-70B-4bits-r4 80.94% 60.00% 76.01% 80.77% 98.88% 96.38% - - -
0.8019 0.6719 0.7219 0.8101 0.9887 0.9474 - - -
@@
Llama-3.1-8B-4bits-r4 16 $\times$ 4 83.6 0.77 $\times$ 4 4.5 0.1
Llama-3.1-8B-8bits-r8 16 $\times$ 4 96.7 0.90 $\times$ 4 9.0 0.1
Llama-3.1-70B-4bits-r4 4 $\times$ 4 184.3 3.50 $\times$ 4 21.3 0.9
Table 5: Finetuning and inference performance on XBRL extraction.
@@
Batch size GPU memory (GB) GPU hours Adapter size (MB) Time (s)
Llama-3.1-8B-4bits-r4 2 $\times$ 4 139.2 0.44 $\times$ 4 4.5 1.9
Llama-3.1-8B-8bits-r8 2 $\times$ 4 152.2 0.48 $\times$ 4 9.0 1.9
@@
  • News with GPT instruction (NWGI) (Liu et al. 2023) comprises samples with seven labels ranging from “strong negative” to “strong positive”. We map the seven labels back to three for simplicity and consistancy with other SA dataset.
@@
The Headline dataset (Sinha and Khandait 2020) categorizes headlines into two classes, ”yes” and ”no”, based on predefined questions.
@@
The NER dataset (Salinas Alvarado, Verspoor, and Baldwin 2015) annotates one entity per sentence, categorized into one of three classes: ”location”, ”person”, and ”organization”
XBRL tagging
The FiNER dataset (Loukas et al. 2022) includes sentences annotated with 139 types of XBRL Tags. We processed the dataset so each question comprises of the sentence and one highlighted entity and the answer includes the correct tag.
XBRL extraction
The XBRL extraction dataset comprises questions and answers derived from XBRL filings from 2019 to 2023 for Dow Jones 30 companies. Each example includes a question, a text segment from an XBRL file containing the answer, and the ground truth generated using an XBRL file extraction library. From this dataset, we selected the following two tasks:
  • XBRL tag extraction: The extraction of a specific XBRL tag from a large XBRL raw text segment given a natural language description of the tag.
  • XBRL value extraction: The extraction of a numeric value from a large XBRL raw text segment given a natural language description of the value.
@@
  • General financial tasks and XBRL tagging: For sentiment analysis, headline classification, and named entity recognition, and XBRL Tagging, single-task fine-tuning was employed.
  • XBRL Extraction: For XBRL tag extraction and value extraction, multi-task fine-tuning was adopted.
@@
  • General financial tasks and XBRL tagging:
@@
Tables 3 summarize the accuracy and weighted F1 scores under different finetuning configurations. Table 4 and 5 displays resource usage and inference performance for NER and XBRL extraction. The finetuned Llama 3.1 8B demonstrates noticeable improvements in accuracy compared to its base model and even surpasses the results of the Llama 3.1 70B base model.
@@
This study demonstrated the effectiveness of Quantized LoRA (QLoRA) for finetuning large language models (LLMs) for many financial tasks, including sentiment analysis, named entity recognition, news headline analysis, and XBRL filings. We finetuned both Llama 3.1 8B and 70B models on commodity GPUs, achieving up to 48% improvements in accuracy compared to the base models on average across all tasks. Notably, these performance gains can be achieved with only four GPUs and less than 20 hours of training times per task, making local finetuning and deployment of customized models a feasible option for institutions and individuals.
In future work, we plan to explore multi-task finetuning in classification tasks and expand our investigation of XBRL-related tasks. This will enable FinLoRA to perform more complex analysis and reasoning tasks, further increasing their utility in the financial domain.
Acknowledgement
Dannong Wang, Daniel Kim, Bo Jin, Xingjian Zhao, Steve Yang and Xiao-Yang Liu Yanglet acknowledge the support from a NSF IUCRC CRAFT center research grant (CRAFT Grant 22017) for this research. The opinions expressed in this publication do not necessarily represent the views of NSF IUCRC CRAFT. Xiao-Yang Liu Yanglet acknowledges the support from Columbia’s SIRS and STAR Program, as well as The Tang Family Fund for Research Innovations in FinTech, Engineering, and Business Operations.
Appendix
Table 6: Finetuning LLMs with QLoRA methods: listing the number of parameters and GPU memory.
Model Parameters GPU Memory Batch = 4 GPU Memory Batch = 8 Model size Percentage
Llama3-8B-16bit (base) 8.03 B - - 16.06 GB -
Llama3-8B-r8-16bit 4.72 M 30.91 GB 30.91 GB 16.08 GB 100.1%
Llama3-8B-r8-8bit 4.72 M 11.41 GB 11.81 GB 8.04 GB 50.1%
Llama3-8B-r8-4bit 4.72 M 8.26 GB 8.65 GB 4.02 GB 25.0%
Llama3-8B-r4-16bit 2.36 M 30.90 GB 30.90 GB 16.07 GB 100.1%
Llama3-8B-r4-8bit 2.36 M 11.40 GB 11.78 GB 8.03 GB 50.0%
Llama3-8B-r4-4bit 2.36 M 8.25 GB 8.61 GB 4.02 GB 25.0%
Llama3-70B-16bit (base) 70.56 B - - 151.53 GB -
Llama3-70B-r8-16bit 22.28 M - - 151.57 GB 100.0%
Llama3-70B-r8-8bit 22.28 M 173.57 GB 258.17 GB 75.79 GB 50.0%
Llama3-70B-r8-4bit 22.28 M 42.78 GB 42.78 GB 37.89 GB 25.0%
Llama3-70B-r4-16bit 11.14 M - - 151.55 GB 100.0%
Llama3-70B-r4-8bit 11.14 M 173.36 GB 258.11 GB 75.70 GB 50.0%
Llama3-70B-r4-4bit 11.14 M 42.73 GB 42.73 GB 37.89 GB 25.0%